# Chemistry-Inspired Diffusion with Non-Diffe-rentiable Guidance

### v1

@@ 1 Introduction
Both explicitly and implicitly guided diffusion require a labeled dataset to train the model or the property predictor. The acquisition of labels, however, can be expensive, time-consuming, and often impracticable [^1] For example, $\text{LD}_{50}$ (Erhirhie et al., 2018) measures the lethal dose of a test substance, which is difficult to calculate from the molecular structure of the substance alone; thus, requiring animal testing.. When only small labeled datasets are available (Power et al., 2022), the model and the property predictor can potentially struggle to generalize beyond seen structures (see Appendix F.2), degrading the performance of the diffusion process when generating novel molecules. Implicitly guided diffusion can help address such challenges. This approach (i) requires no labels to train the model and (ii) can replace the property predictor with domain knowledge from quantum chemistry, fulfilled by quantum chemistry software such as xTB (Bannwarth et al., 2019; 2021) and Gaussian (Frisch et al., 2016). Such software can act as an expert oracle to create labeled datasets; thereby, avoiding the extrapolation shortcomings of neural networks.
Despite following certain computation procedures, quantum chemistry is by nature a non-differentiable oracle [^2] Here it refers to the non-differentiability across the chemical space between different molecules. that can not be backpropagated with neural networks, as those procedures usually involve proprietary algorithms or extremely computationally expensive to perform, which prevents the integration of quantum chemistry to guide the diffusion models. In this work, we aim to implicitly guide the diffusion process with a non-differentiable oracle, and employ zeroth-order optimization methods (Nesterov & Spokoiny, 2017; Malladi et al., 2023) to estimate the necessary guidance gradients. We refer to our method as ChemGuide, a method that leverages quantum chemistry to provide more accurate diffusion guidance. Specifically, given the oracle xTB with the GFN2-xTB method (Bannwarth et al., 2019) that conducts quantum chemistry calculation of potential energies and atomic forces, we first use ChemGuide to improve the stability of the generated molecular geometry (see Fig. 1), by guiding the diffusion process towards sampling from a distribution where the net force on the atoms is zero. Further, we formulate diffusion guidance on both molecular property and stability as a bilevel optimization (Colson et al., 2007) problem, to demonstrate that ChemGuide is compatible with both explicitly and implicitly guided diffusion. We summarize our contributions as follows:
  • We propose ChemGuide, which leverages zeroth-order optimization techniques to introduce a non-differentiable chemistry oracle as diffusion guidance (Section 3). To the best of our knowledge, this is the first work of its kind.
@@ 2 Preliminary
However, the exact mathematical potential energy evaluation, i.e., the solution to the Schrödinger equation, is still a “black box” (Cao et al., ). Over the years, different levels of theories and methods have been developed to evaluate potential energies, such as force field (FF), semi-empirical methods (e.g., xTB), and density functional theory (DFT) methods (e.g., B3LYP/6-31G(2df,p)), listed in an order of increased accuracy and cost. After trading-off between accuracy and efficiency within a feasible computation cost, we selected GFN2-xTB, a more recent and advanced semi-empirical method (Bannwarth et al., 2019), to calculate the forces of the generated molecular geometry in the diffusion process. More details about GFN2-xTB can be found in Appendix C.
@@
Stable molecular geometry is favored for molecule generation. A novel but unstable molecular geometry may not be physically achieved at all and thus carries no substantial meaning. To optimize stability, one can minimize the potential energy of the molecule by iteratively minimizing the forces on the atoms (Eq. 5), motivating our choice of a 3D diffusion model. More details and discussions are provided in Appendix D.
3 Methodology
3.1 Guidance From Neural Regressor
@@ 3.2 Guidance From a Non-Differentiable Oracle
^$\hat{\nabla}_{{\bm{z}}_{t}}\log p_{\phi}(y|{\bm{z}}_{t})$ $\propto-\nabla_{\mathcal{F}({\bm{z}}_{t})}\mathcal{L}(y,\mathcal{F}({\bm{z}}_{t}))\frac{\mathcal{F}([{\bm{z}}_{\mathbf{x},t}+\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}])-\mathcal{F}([\mathbf{z}_{\mathbf{x},t}+\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}])}{2\zeta}\mathbf{U}$ (15)
where $\zeta$ is a small perturbation scale (e.g., $10^{-6}$). The perturbed representations $[{\bm{z}}_{\mathbf{x},t}\pm\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}]$ obey zero central gravity required by $t_{0}$, as $\mathbb{E}[\zeta\mathbf{U}]=0$. Note that, unlike neural regressor guidance, we only add guidance to the positions (i.e. ${\bm{z}}_{\mathbf{x},t}$) and apply no gradient to the atom types (i.e. ${\bm{z}}_{\mathbf{h},t}$). We do this because the force definition (Eq. 5) is only physically grounded when the set of atoms stays constant (i.e. no matter/mass is created from or reduced to void). In other words, since energy is a relative concept, any gradient that arises from atom-type change has no physical interpretation.
@@ 3.3 Combine Guidance from Neural Network and Non-Differentiable Oracle
     $g_{t-1}\propto-\nabla_{\mathcal{F}({\bm{z}}_{0}^{\prime})}\|\mathcal{F}({\bm{z}}_{0}^{\prime})\|^{2}\cdot\frac{1}{2\zeta}\left(\mathcal{F}([{\bm{z}}_{\mathbf{x},0}^{\prime}+\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},0}^{\prime}])-\mathcal{F}([{\bm{z}}_{\mathbf{x},0}^{\prime}-\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},0}^{\prime}])\right)\mathbf{U}$ (Eq. 15)
@@
where $y$ is the target property value (e.g., we hope the static polarizability $\alpha\rightarrow y$ after optimization). As detailed in Algorithm 1, we first optimize molecular property with $f_{\eta}$, and obtain the optimal atom positions ${\bm{z}}_{x,t}^{*}$ and features ${\bm{z}}_{h,t}^{*}$ (Eq. 17); we then optimize stability with $\mathcal{F}$ by fixing $\mathbf{z}_{h,t}^{*}$ and initializing ${\bm{z}}_{x,t}$ (Eq. 16) as ${\bm{z}}_{x,t}^{*}$. The chemistry intuition of our design is detailed in Appendix E.
@@ 4.1 Experiment Setting
We add guidance to the last 400 of the 1000 diffusion steps (Han et al., 2024), and calculate the change percentage between our method and the GeoLDM/C-GeoLDM baseline. The implementation details are provided in Appendix A.
@@ 4.2 Non-Differentiable Oracle Guidance on Force
Fig. 2 shows the plot of the force RMS and energy change histogram and distribution of the 500 generated molecules from EDM, GeoLDM, and ChemGuide on QM9 and GEOM. Our method is more concentrated around 0 and less spread to higher regions. Also, note that both EDM and GeoLDM generate outlier molecules with significantly large force and energy change, while no outlier is generated by ChemGuide. For example, in Fig. 2(a), an outlier with force RMS 0.155 is generated by GeoLDM, which is $\sim$15 times higher than the force RMS where the molecules are concentrated. It demonstrates that our method performs better both on average and molecule-wise, and we provide detailed plots of various scales in Fig. 10 in Appendix H.
@@ 4.3.1 Unconditional Diffusion with noisy neural guidance
We present the results of noisy guidance (Section 3.1) in Table 4 and provide further details in Table 11 in Appendix I. For all the properties $\alpha$, $\mu$, $C_{v}$, $\epsilon_{HOMO}$, $\epsilon_{LUMO}$, and $\Delta\epsilon$, noisy neural guidance outperforms unconditional GeoLDM by 57.92%, 37.24%, 77.09%, 23.18%, 36.87%, and 37.88% respectively, which verifies its effectiveness. Furthermore, GeoLDM with noisy neural guidance yields comparable performance with C-GeoLDM and C-EDM, provided that conditional models are explicitly trained for property optimization; surprisingly, we use unconditional GeoLDM as the backbone and outperforms C-GeoLDM in $\alpha$, $\mu$, and $C_{v}$. To understand the details of guided generation, we plot the MAE of $C_{v}$ on sampled molecules at each guidance step in Fig. 3. Compared to GeoLDM which oscillates in the first 200 steps and converges slowly to a high MAE, the trajectory of noisy neural guidance drops fast in the first 50 steps and the amplitude of the oscillation is less intense, which suggests the guidance smooths the MAE change and quickly steers the model toward better properties. The property MAE trajectories are detailed in Fig. 11 in Appendix I.
@@ 4.3.2 Unconditional Diffusion with clean neural guidance
As presented in Table 4, clean neural guidance outperforms unconditional GeoLDM on all properties and beats C-GeoLDM in $\mu$, $C_{v}$, $\epsilon_{\text{HOMO}}$, and $\Delta\epsilon$ by $81.68\%$, $67.53\%$, $3.06\%$, and $34.84\%$. Interestingly, it seems that noisy and clean guidance complement each other on different properties (e.g., $\alpha,\Delta\epsilon$). In Fig. 3, clean neural guidance shows a similar oscillation trend in the first 200 guidance steps compared to GeoLDM, then exhibits a sharp drop in the last 50 steps and achieves the lowest MAE, which is also observed on other properties shown in Fig. 11 in Appendix I. We hypothesize this is due to that molecules are less noisy as denoising steps accumulate; thus, guidance optimization in the clean space manifests its strongest power when ${\bm{z}}_{t}$ is less dependent on the one-step estimation function $t_{0}(\cdot)$ in later steps, where it is closer and more similar to ${\bm{z}}_{0}$ (Eq. 8).
@@ 4.4.1 Explicitly Guided Diffusion with ChemGuide
We employ C-GeoLDM with ChemGuide, which resembles the bilevel optimization framework as C-GeoLDM automatically optimizes molecular property during the denoising process. Compared to C-GeoLDM alone, we expect our method to reduce the force while not hurting MAE. We present the results in Fig. 5 and Table 5 (details are provided in Table 12 and Table 13 in Appendix J). For the 6 properties ($\alpha$, $\mu$, $C_{v}$, $\epsilon_{HOMO}$, $\epsilon_{LUMO}$, $\Delta\epsilon$), our method reduces the force RMS by 9.21%, 3.33%, 8.75%, 6.69%, 1.27%, and 10.19%, and for the first 5, MAEs decrease 4.54%, 9.49%, 2.62%, 0.09%, and 14.34%. The results suggest it is feasible and effective to combine explicitly guided diffusion (e.g., C-GeoLDM) with ChemGuide, which goes beyond our expectation as both force RMS and property MAE are reduced.
@@ 4.4.2 Implicitly Guided Diffusion with ChemGuide
We apply two aforementioned neural network guidance approaches (noisy/clean guidance, Section 3.1) for the property optimization step in our bilevel optimization framework (Section 3.3). We expect that both the property MAE and the force metrics are better than conditional GeoLDM.
@@
As shown in Table 5, our method outperforms C-GeoLDM in $\alpha$, $\mu$, and $C_{v}$ by 2.86%, 37.43%, and 2.87% in MAE. It is also observed in Fig. 5 that, for the six properties, our method achieves lower force RMS than C-GeoLDM by 1.94%, 4.57%, 8.97%, 10.66%, 16.26%, and 7.18%, with smaller energy change compared to baselines. Details of MAE and force metrics are provided in Table. 13 and Table. 14 in Appendix J. Combining property optimization with force optimization, noisy guidance with ChemGuide reduces MAE and improves the stability of generated molecules, which suggests that the combined guidance prompts the diffusion model towards generating stabler molecules with better properties.
To further understand how our bilevel framework optimizes property, we provide the MAE trajectory of $C_{v}$ in Fig. 5. The trajectories of our noisy neural guidance method (Fig.3) and bilevel optimization behave similarly, such that the MAE drops fast in the first 50 steps and oscillates less compared to GeoLDM. However, bilevel optimization is slightly worse than the noisy neural guidance method (i.e. it converges closer to C-GeoLDM than noisy guidance). We suspect it is because guided by gradients from the neural regressor and then ChemGuide, the generation process leans less toward a lower MAE, as it keeps a balance between property and stability optimization. Details and explanations of the MAE trajectories of each property can be found in Fig. 12 in Appendix J.
@@
We present the details of this section in Appendix G.

### v2

@@ Abstract
[…]
https://github.com/A-Chicharito-S/ChemGuide
11footnotetext: Equal contribution.  Correspondence to: Barnabás Póczos and Newell Washburn.
@@ 1 Introduction
Both explicitly and implicitly guided diffusion require a labeled dataset to train the model or the property predictor. The acquisition of labels, however, can be expensive, time-consuming, and often impracticable [^1] For example, $\text{LD}_{50}$ (Erhirhie et al., 2018) measures the lethal dose of a test substance, which is difficult to calculate from the molecular structure of the substance alone; thus, requiring animal testing.. When only small labeled datasets are available (Power et al., 2022), the model and the property predictor can potentially struggle to generalize beyond seen structures (see Appendix H.4), degrading the performance of the diffusion process when generating novel molecules. Implicitly guided diffusion can help address such challenges. This approach (i) requires no labels to train the model and (ii) can replace the property predictor with domain knowledge from quantum chemistry, fulfilled by quantum chemistry software such as xTB (Bannwarth et al., 2019; 2021) and Gaussian (Frisch et al., 2016). Such software can act as an expert oracle to create labeled datasets; thereby, avoiding the extrapolation shortcomings of neural networks.
Despite following certain computation procedures, quantum chemistry is by nature a non-differentiable oracle that can not be backpropagated with neural networks, as those procedures usually involve proprietary algorithms or extremely computationally expensive to perform, which prevents the integration of quantum chemistry to guide the diffusion models. In this work, we aim to implicitly guide the diffusion process with a non-differentiable oracle, and employ zeroth-order optimization methods (Nesterov & Spokoiny, 2017; Malladi et al., 2023) to estimate the necessary guidance gradients. We refer to our method as ChemGuide, a method that leverages quantum chemistry to provide more accurate diffusion guidance. Specifically, given the oracle xTB with the GFN2-xTB method (Bannwarth et al., 2019) that conducts quantum chemistry calculation of potential energies and atomic forces, we first use ChemGuide to improve the stability of the generated molecular geometry (see Fig. 1), by guiding the diffusion process towards sampling from a distribution where the net force on the atoms is zero. Further, we formulate diffusion guidance on both molecular property and stability as a bilevel optimization (Colson et al., 2007) problem, to demonstrate that ChemGuide is compatible with both explicitly and implicitly guided diffusion. We summarize our contributions as follows:
  • We propose ChemGuide, which leverages zeroth-order optimization techniques to introduce a non-differentiable chemistry oracle as diffusion guidance (Section 3).
@@ 2 Preliminary
However, the exact mathematical potential energy evaluation, i.e., the solution to the Schrödinger equation, is still a “black box” (Cao et al., ). Over the years, different levels of theories and methods have been developed to evaluate potential energies, such as force field (FF), semi-empirical methods (e.g., xTB), and density functional theory (DFT) methods (e.g., B3LYP/6-31G(2df,p)), listed in an order of increased accuracy and cost. After trading-off between accuracy and efficiency within a feasible computation cost, we selected GFN2-xTB, a more recent and advanced semi-empirical method (Bannwarth et al., 2019), to calculate the forces of the generated molecular geometry in the diffusion process. More details about GFN2-xTB can be found in Appendix D.
@@
Stable molecular geometry is favored for molecule generation. A novel but unstable molecular geometry may not be physically achieved at all and thus carries no substantial meaning. To optimize stability, one can minimize the potential energy of the molecule by iteratively minimizing the forces on the atoms (Eq. 5), motivating our choice of a 3D diffusion model. More details and discussions are provided in Appendix E.
@@ 3.1 Guidance From a Non-Differentiable Oracle
^$\hat{\nabla}_{{\bm{z}}_{t}}\log p_{\phi}(y|{\bm{z}}_{t})$ $\propto-\nabla_{\mathcal{F}({\bm{z}}_{t})}\mathcal{L}(y,\mathcal{F}({\bm{z}}_{t}))\frac{\mathcal{F}([{\bm{z}}_{\mathbf{x},t}+\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}])-\mathcal{F}([\mathbf{z}_{\mathbf{x},t}-\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}])}{2\zeta}\mathbf{U}$ (15)
where $\zeta$ is a small perturbation scale (e.g., $10^{-6}$). The perturbed representations $[{\bm{z}}_{\mathbf{x},t}\pm\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}]$ obey zero central gravity required by $t_{0}$, as $\mathbb{E}[\zeta\mathbf{U}]=0$. Note that, unlike neural regressor guidance, we only add guidance to the positions (i.e. ${\bm{z}}_{\mathbf{x},t}$) and apply no gradient to the atom types (i.e. ${\bm{z}}_{\mathbf{h},t}$). We do this because the force definition (Eq. 5) is only physically grounded when the set of atoms stays constant (i.e. no matter/mass is created from or reduced to void). In other words, since energy is a relative concept, any gradient that arises from atom-type change has no physical interpretation. Further discussions on the choice of SPSA for gradient estimation can be found in Appendix F.
@@ 3.2 Combine Guidance from Neural Network and Non-Differentiable Oracle
     $g_{t-1}\propto-\nabla_{\mathcal{F}({\bm{z}}_{t}^{\prime})}\|\mathcal{F}({\bm{z}}_{t}^{\prime})\|^{2}\cdot\frac{1}{2\zeta}\left(\mathcal{F}([{\bm{z}}_{\mathbf{x},t}^{\prime}+\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}^{\prime}])-\mathcal{F}([{\bm{z}}_{\mathbf{x},t}^{\prime}-\zeta\mathbf{U},{\bm{z}}_{\mathbf{h},t}^{\prime}])\right)\mathbf{U}$ (Eq. 15)
@@
     ${\bm{z}}_{t-1}\leftarrow\mathcal{N}(\mu_{t-1}^{\prime}+s\cdot\sigma_{t-1}^{2}g_{t-1},\Sigma_{t-1}^{\prime})$ (Eq. 6)
@@
where $y$ is the target property value (e.g., we hope the static polarizability $\alpha\rightarrow y$ after optimization). As detailed in Algorithm 2, we first optimize molecular property with $f_{\eta}$, and obtain the optimal atom positions ${\bm{z}}_{x,t}^{*}$ and features ${\bm{z}}_{h,t}^{*}$ (Eq. 17); we then optimize stability with $\mathcal{F}$ by fixing $\mathbf{z}_{h,t}^{*}$ and initializing ${\bm{z}}_{x,t}$ (Eq. 16) as ${\bm{z}}_{x,t}^{*}$. The chemistry intuition of our design is detailed in Appendix G.
@@ 4.1 Experiment Setting
We add guidance to the last 400 of the 1000 diffusion steps (Han et al., 2024), and calculate the change percentage between our method and the GeoLDM/C-GeoLDM baseline. The implementation details are provided in Appendix B.
@@ 4.2 Non-Differentiable Oracle Guidance on Force
Fig. 2 shows the plot of the force RMS and energy change histogram and distribution of the 500 generated molecules from EDM, GeoLDM, and ChemGuide on QM9 and GEOM. Our method is more concentrated around 0 and less spread to higher regions. Also, note that both EDM and GeoLDM generate outlier molecules with significantly large force and energy change, while no outlier is generated by ChemGuide. For example, in Fig. 2(a), an outlier with force RMS 0.155 is generated by GeoLDM, which is $\sim$15 times higher than the force RMS where the molecules are concentrated. It demonstrates that our method performs better both on average and molecule-wise, and we provide detailed plots of various scales in Fig. 11 in Appendix K.
@@ 4.3.1 Unconditional Diffusion with noisy neural guidance
We present the results of noisy guidance (Section 2) in Table 4 and provide further details in Table 16 in Appendix L. For all the properties $\alpha$, $\mu$, $C_{v}$, $\epsilon_{HOMO}$, $\epsilon_{LUMO}$, and $\Delta\epsilon$, noisy neural guidance outperforms unconditional GeoLDM by 57.92%, 37.24%, 77.09%, 23.18%, 36.87%, and 37.88% respectively, which verifies its effectiveness. Furthermore, GeoLDM with noisy neural guidance yields comparable performance with C-GeoLDM and C-EDM, provided that conditional models are explicitly trained for property optimization; surprisingly, we use unconditional GeoLDM as the backbone and outperforms C-GeoLDM in $\alpha$, $\mu$, and $C_{v}$. To understand the details of guided generation, we plot the MAE of $C_{v}$ on sampled molecules at each guidance step in Fig. 3. Compared to GeoLDM which oscillates in the first 200 steps and converges slowly to a high MAE, the trajectory of noisy neural guidance drops fast in the first 50 steps and the amplitude of the oscillation is less intense, which suggests the guidance smooths the MAE change and quickly steers the model toward better properties. The property MAE trajectories are detailed in Fig. 12 in Appendix L.
@@ 4.3.2 Unconditional Diffusion with clean neural guidance
As presented in Table 4, clean neural guidance outperforms unconditional GeoLDM on all properties and beats C-GeoLDM in $\mu$, $C_{v}$, $\epsilon_{\text{HOMO}}$, and $\Delta\epsilon$ by $81.68\%$, $67.53\%$, $3.06\%$, and $34.84\%$. Interestingly, it seems that noisy and clean guidance complement each other on different properties (e.g., $\alpha,\Delta\epsilon$). In Fig. 3, clean neural guidance shows a similar oscillation trend in the first 200 guidance steps compared to GeoLDM, then exhibits a sharp drop in the last 50 steps and achieves the lowest MAE, which is also observed on other properties shown in Fig. 12 in Appendix L. We hypothesize this is due to that molecules are less noisy as denoising steps accumulate; thus, guidance optimization in the clean space manifests its strongest power when ${\bm{z}}_{t}$ is less dependent on the one-step estimation function $t_{0}(\cdot)$ in later steps, where it is closer and more similar to ${\bm{z}}_{0}$ (Eq. 8).
@@ 4.4.1 Explicitly Guided Diffusion with ChemGuide
We employ C-GeoLDM with ChemGuide, which resembles the bilevel optimization framework as C-GeoLDM automatically optimizes molecular property during the denoising process. Compared to C-GeoLDM alone, we expect our method to reduce the force while not hurting MAE. We present the results in Fig. 5 and Table 5 (details are provided in Table 17 and Table 18 in Appendix M). For the 6 properties ($\alpha$, $\mu$, $C_{v}$, $\epsilon_{HOMO}$, $\epsilon_{LUMO}$, $\Delta\epsilon$), our method reduces the force RMS by 9.21%, 3.33%, 8.75%, 6.69%, 1.27%, and 10.19%, and for the first 5, MAEs decrease 4.54%, 9.49%, 2.62%, 0.09%, and 14.34%. The results suggest it is feasible and effective to combine explicitly guided diffusion (e.g., C-GeoLDM) with ChemGuide, which goes beyond our expectation as both force RMS and property MAE are reduced.
@@ 4.4.2 Implicitly Guided Diffusion with ChemGuide
We apply two aforementioned neural network guidance approaches (noisy/clean guidance, Section 2) for the property optimization step in our bilevel optimization framework (Section 3.2). We expect that both the property MAE and the force metrics are better than conditional GeoLDM.
@@
As shown in Table 5, our method outperforms C-GeoLDM in $\alpha$, $\mu$, and $C_{v}$ by 2.86%, 37.43%, and 2.87% in MAE. It is also observed in Fig. 5 that, for the six properties, our method achieves lower force RMS than C-GeoLDM by 1.94%, 4.57%, 8.97%, 10.66%, 16.26%, and 7.18%, with smaller energy change compared to baselines. Details of MAE and force metrics are provided in Table. 18 and Table. 19 in Appendix M. Combining property optimization with force optimization, noisy guidance with ChemGuide reduces MAE and improves the stability of generated molecules, which suggests that the combined guidance prompts the diffusion model towards generating stabler molecules with better properties.
To further understand how our bilevel framework optimizes property, we provide the MAE trajectory of $C_{v}$ in Fig. 5. The trajectories of our noisy neural guidance method (Fig.3) and bilevel optimization behave similarly, such that the MAE drops fast in the first 50 steps and oscillates less compared to GeoLDM. However, bilevel optimization is slightly worse than the noisy neural guidance method (i.e. it converges closer to C-GeoLDM than noisy guidance). We suspect it is because guided by gradients from the neural regressor and then ChemGuide, the generation process leans less toward a lower MAE, as it keeps a balance between property and stability optimization. Details and explanations of the MAE trajectories of each property can be found in Fig. 13 in Appendix M.
@@
We present the details of this section in Appendix J.